# A Survey of Text Classification Under Class Distribution Shift

### v2

@@ 8 Limitations
Appendix A Text Datasets for Class Distribution Shift
@@
AG News Zhang et al. (2015) Topic categorization 127k 4
DBPedia Lehmann et al. (2014) Topic categorization 630k 14
20 Newsgroup Lang (1995) Topic categorization 18k 20
Yahoo! Answers Zhang et al. (2015) Topic categorization 1.4m 10
SST-2 Socher et al. (2013) Polarity classification 69k 2
IMDb Maas et al. (2011) Polarity classification 50k 2
Amazon Product Reviews Chen et al. (2013) Topic modeling 50k 50
TMNtitle Gupta et al. (2020) Topic modeling 32.6k 7
R21578title Gupta et al. (2020) Topic modeling 10.8k 90
Table 2: List of datasets that are commonly used in text classification under class distribution shift. For the selection of datasets, we report some of their basic statistics, as well as the tasks where they are employed.
@@ 20 Newsgroup Lang (1995) Topic categorization 18k 20
A possible methodology for zero-shot testing is proposed by Yin et al. (2019), alongside a few suitable datasets. For topic categorization, the authors propose the Yahoo! Answers Zhang et al. (2015) dataset. This dataset is reorganized into dev and test splits, both containing all 10 labels. Dev consists of 6k instances per label, and test has 10k instances per label. For the labels-partially-unseen scenario, two variants of the dataset are created, with different partitions of seen/unseen classes, both balanced. For labels-fully-unseen there is no training set. Sanh et al. (2022) employ AG News Zhang et al. (2015) and DBPedia Lehmann et al. (2014) for zero-shot topic categorization, and IMDb Maas et al. (2011) for polarity classification. There is no training set for the target task. Similarly, Meng et al. (2020) uses the AG News, DBPedia, IMDb and Amazon Product Reviews Chen et al. (2013), with all the labels removed. In general, we observe that the preferred datasets are large scale.
Text datasets for open-set classification.
@@
The common way to test an open-set solution is to reorganize a dataset designed for supervised text classification. To this end, some classes are reserved to play the role of the open space and are not used during training. While this scenario may seem similar to the zero-shot labels-partially-unseen setup, the labels for the reserved classes are not used. During testing, the system must simply detect them as belonging to the open space, without classifying them into classes. Many works perform multiple experiments, varying the KKCs-to-UUCs ratio between 20% and 100%. It is worth noting that a percentage of 100% known classes is not a closed-set problem, because the model can still classify instances as open space.
@@
Kim et al. (2022) report experiments on AG News and DBPedia. The KKCs-to-UUCs ratio is chosen among three values: 25%, 50% and 75%. For AG News, all the combinations of known / unknown classes having at least one known and one unknown class are compared. Some studies Chen et al. (2023, 2024) conduct experiments on three datasets, namely AG News, DBPedia and Yahoo! Answers, reserving two, four or six classes to play the role of UUCs. Other papers Walkowiak et al. (2019b, 2018) use custom datasets, specifically tailored for the open-set task. Walkowiak et al. (2019b) are the only ones who use foreign language data (Polish).
Text datasets for continual learning.
@@
Gupta et al. (2020) report experiments on Amazon Product Reviews Chen et al. (2013), 20NSshort, TMNtitle and R21578title datasets. They use these datasets as future tasks in a continual learning framework for topic modeling, testing the ability of models to learn from short texts, retain knowledge, and adapt to new domains without catastrophic forgetting. The Amazon Product Reviews dataset is preprocessed according to the methodology described by Chen et al. (2013). The preprocessing includes sentence detection, lemmatization, and POS tagging, as well as the removal of punctuation, stop words, and rare terms. To avoid excessive thematic overlap, the domain name of each collection is removed. 20NSshort is a subset of 20 Newsgroups Lang (1995), containing only documents with fewer than 20 words, making topic inference challenging due to sparse content. Similarly, R21578title is derived from Reuters-21578 Lewis (1987), retaining only article titles, further limiting contextual information. TMNtitle comes from Tag My News (TMN) Vitale et al. (2012) and includes only news headlines, reflecting real-world short-text classification tasks.

### v3

@@ Acknowledgments
Appendix A Appendix
In the supplementary, we discuss related work on continual learning from text and how this research direction is useful in text classification under class distribution shift (Section A.1). We also review the datasets that are commonly used in zero-shot and open-set learning from text (Section A.2).
A.1 Continual Learning from Text
Formal definition. Continual learning (CL) is a paradigm in machine learning focused on enabling systems to continuously learn from a stream of data over time. Unlike traditional approaches that assume a static data distribution and rely on fixed datasets, CL addresses challenges such as retaining knowledge from previous tasks (avoiding catastrophic forgetting (Kirkpatrick et al., 2017)) and adapting to new tasks without extensive retraining (Parisi et al., 2019). CL can be defined as the process of training a model on a sequence of tasks, where each task $t$ can be characterized by a distinct data distribution. At any point in time, the model receives a batch of training samples $D_{t,b}=\{(x,y)\mid x\in\mathcal{X}_{t,b},y\in\mathcal{Y}_{t,b}\}$, where $\mathcal{X}_{t,b}$ represents a subset of input samples and $\mathcal{Y}_{t,b}$ are the corresponding (task-specific) labels. The task identity is given by $t\in\{1,$$\infty\}$, while the batch index is represented by $b\in\{1,$$\infty\}.$ Theoretically, the continual learning process continues indefinitely during operation, hence the indexes $t$ and $b$ can grow endlessly. A task $t$ is formally defined by its data distribution $P(\mathcal{X}_{t},\mathcal{Y}_{t})$ (Wang et al., 2024).
Trends and approaches. CL is a relevant framework for open-set learning, as it can address challenges related to the open-set scenarios, especially in tasks such as text classification by topic when the topic distribution changes. These challenges include handling new classes and adapting to an evolving data distribution. CL can help to discover and learn new topics during inference. Both Chen and Liu (2014) and Gupta et al. (2020) propose frameworks that enable models to accumulate knowledge over time, leveraging past knowledge to improve topic extraction, while mitigating catastrophic forgetting. These approaches integrate lifelong knowledge retention and transfer, ensuring that previously learned topics inform future learning tasks. By incorporating prior knowledge, these models enhance topic coherence, particularly in sparse data scenarios, where limited labeled data can affect the quality of extracted topics. Chen and Liu (2014) introduce the Lifelong Topic Model (LTM), which mines prior knowledge from past document collections and integrates it into new topic modeling tasks through probabilistic inference. Similarly, Gupta et al. (2020) propose the Lifelong Neural Topic Modeling (LNTM) framework, which relies on neural networks and selective data augmentation to balance stability and plasticity. LTM is designed for multi-domain topic extraction, identifying topic structures across different domains, while LNTM is more suitable for continuous document streams, maintaining topic stability, while adapting to novel information. Moreover, LTM employs a knowledge-based topic modeling approach that refines topic quality through extracted knowledge sets, whereas LNTM integrates topic regularization and selective data augmentation to enhance adaptation, while preventing excessive forgetting.
Open-world continual learning. Open-world continual learning is a learning paradigm in which a model incrementally acquires new tasks and, at the same time, detects and integrates previously unseen classes. In their recent work, Li et al. (2025a) propose a framework called HoliTrans, which transfers knowledge from both known and unknown class samples. HoliTrans uses Nonlinear Random Projection to create clearer representations that help to separate new samples from previously learned classes. The method also relies on Distribution-Aware Prototypes, which dynamically adapt as new classes appear. These components help to better distinguish between new and previously seen classes, particularly when unknown classes repeatedly occur.
A.2 Text Datasets for Class Distribution Shift
@@
AG News (Zhang et al., 2015) Topic categorization 127k 4
DBPedia (Lehmann et al., 2014) Topic categorization 630k 14
20 Newsgroup (Lang, 1995) Topic categorization 18k 20
Yahoo! Answers (Zhang et al., 2015) Topic categorization 1.4m 10
SST-2 (Socher et al., 2013) Polarity classification 69k 2
IMDb (Maas et al., 2011) Polarity classification 50k 2
Amazon Product Reviews (Chen et al., 2013) Topic modeling 50k 50
TMNtitle (Gupta et al., 2020) Topic modeling 32.6k 7
R21578title (Gupta et al., 2020) Topic modeling 10.8k 90
Table 3: List of datasets that are commonly used in text classification under class distribution shift. For the selection of datasets, we report some of their basic statistics, as well as the tasks where they are commonly employed.
@@ 20 Newsgroup (Lang, 1995) Topic categorization 18k 20
Text datasets for zero-shot classification. A possible methodology for zero-shot testing is proposed by Yin et al. (2019), alongside a few suitable datasets. For topic categorization, the authors propose the Yahoo! Answers (Zhang et al., 2015) dataset. This dataset is reorganized into dev and test splits, both containing all 10 labels. The dev split consists of 6k instances per label, while the test split has 10k instances per label. For the labels-partially-unseen scenario, two variants of the dataset are created, with different partitions of seen/unseen classes, both balanced. For labels-fully-unseen there is no training set. Sanh et al. (2022) employ AG News (Zhang et al., 2015) and DBPedia (Lehmann et al., 2014) for zero-shot topic categorization, and IMDb (Maas et al., 2011) for polarity classification. There is no training set for the target task. Similarly, Meng et al. (2020) use the AG News, DBPedia, IMDb and Amazon Product Reviews (Chen et al., 2013) datasets, with all the labels removed. In general, we observe that the preferred datasets are large scale.
@@
Text datasets for open-set classification. The common way to test an open-set solution is to reorganize a dataset designed for supervised text classification. To this end, some classes are reserved to play the role of the open space and are not used during training. While this scenario may seem similar to the zero-shot labels-partially-unseen setup, the labels for the reserved classes are not used. During testing, the system must simply detect them as belonging to the open space, without classifying them into classes. Many works perform multiple experiments, varying the KKCs-to-UUCs ratio between 20% and 100%. It is worth noting that a percentage of 100% known classes is not a closed-set problem, because the model can still classify instances as open space.
@@
Kim et al. (2022) report experiments on AG News and DBPedia. The KKCs-to-UUCs ratio is chosen among three values: 25%, 50% and 75%. For AG News, all the combinations of known / unknown classes having at least one known and one unknown class are compared. Some studies (Chen et al., 2023, 2024) conduct experiments on three datasets, namely AG News, DBPedia and Yahoo! Answers, reserving two, four or six classes to play the role of UUCs. Other papers (Walkowiak et al., 2018, 2019b) use custom datasets, specifically tailored for the open-set task. Walkowiak et al. (2019b) are the only ones who use data in a foreign language, namely Polish.
@@
Text datasets for continual learning. Gupta et al. (2020) report experiments on Amazon Product Reviews (Chen et al., 2013), 20NSshort, TMNtitle and R21578title datasets. They use these datasets as future tasks in a continual learning framework for topic modeling, testing the ability of models to learn from short texts, retain knowledge, and adapt to new domains without catastrophic forgetting. The Amazon Product Reviews dataset is preprocessed according to the methodology described by Chen et al. (2013). The preprocessing includes sentence detection, lemmatization, and POS tagging, as well as the removal of punctuation, stop words, and rare terms. To avoid excessive thematic overlap, the domain name of each collection is removed. 20NSshort is a subset of 20 Newsgroups (Lang, 1995), containing only documents with fewer than 20 words, making topic inference challenging due to sparse content. Similarly, R21578title is derived from Reuters-21578 (Lewis, 1987), retaining only article titles, further limiting contextual information. TMNtitle comes from Tag My News (TMN) (Vitale et al., 2012) and includes only news headlines, reflecting real-world short-text classification tasks.
@@
General remark. We find that datasets used in literature do not always reflect real-world open-set scenarios. This is mostly due to the fact that nearly all datasets were introduced more than 10 years ago. We thus believe that collecting new resources that cover realistic open-set setups is an important avenue for future research.